Pew research surveyed 5006 US adults to ask their opinion about the state of jobs n the US in 2016. They asked what employment sector respondents worked in and whether their job gave them a sense of identity or whether it was “just what they do for a living”. This table summarizes their responses:Private company Nonprofit Government Self-employed Sense of Identit 42% 65% 67% 62% Just what I do for a livin 55% 34% 32% 36%

Which of the following can you tell from this table? Give the value if you can find it.

  1. The percent of respondents who both consider their job just what they do for a living and are self-employed.
  2. The percent of non-profit employees who get a sense of identity from their job.
  3. The percent of those who find their work “Just what they do for a living” who are Private company employees.
  4. The percent of all respondents who get a sense of identity from their job.
Answer #1

Answer

(A) In this case, we have to find the intersection point of "just what they do for a living" and "self employed". Only 36% people fall under this category which are just what they do for a living and are self employed.

(b) In this case, we have to find the intersection point of non-profit employees and sense of identity. Only 65% people fall under this category.

(C) In this case, wehave to find the intersection point of private company and just what they do for a living. Only 34% people are there who are in private company and just what I do for a living.

(d) In this case, we have to find the average % of private company, non-profit, government and self-employed under the category of "sense of identity"

So, average = (sum of all % under "sense of identity")/(total number of all % under "sense of identity") = (42+65+67+62)/4

this gives us

Average = 236/4= 59%

So, we can say that 59 percent of all respondents who get a sense of identity from their job.